# Path Gradients after Flow Matching

### v2

@@ 1 Introduction
  • We investigate Path Gradients impact on properties trained during Flow Matching, namely flow trajectory length and the MSE Flow Matching loss, and show that these are mostly unaffected by the fine-tuning.
@@ 2.1 Boltzmann Generators
Boltzmann Generators Noé et al. (2019) aim to asymptotically generate unbiased samples from the equilibrium Boltzmann distribution
@@ 2.4 Path Gradients
Roeder et al. (2017) observed, that the martingale score term vanishes in expectation, but has a non-zero variance. Vaitl et al. (2022a) showed that its variance is $\frac{1}{N}I_{\theta}$ for a batch of size $N$ and the Fisher Information $I_{\theta}$ of the variational distribution $q_{\theta}$. The low variance close to the optimum allows the gradient estimator to "stick-the-landing" (Roeder et al., 2017), i.e. having zero variance at the optimum, making it an ideal tool for finetuning. Vaitl et al. (2024) showed that Path Gradient estimators incorporate additional information about the derivative of both target and variational distribution, opposed to the reparametrized gradient gradient estimator (Kingma & Welling, 2014; Mohamed et al., 2020) (see Section˜A.11.1 for a detailed explanation). They hypothesized that this leads to less overfitting and a generally better fit to the target density (see Vaitl (2024)). These theoretical aspects make Path Gradients a promising suitor for fine-tuning Boltzmann Generators.
@@
While the naive calculation of Path Gradients requires significantly more compute and memory than Flow Matching, we take several steps to obtain constant memory and speed up computation in the next section. As we show experimentally, fine-tuning with Path Gradient for a few epochs significantly improves the performance for Continuous Normalizing Flows compared to only using Flow Matching.
@@ 3 Path Gradients and Flow Matching
$\frac{d}{dt}\frac{\partial\log p_{t,\theta}(x_{t})}{\partial x_{t}}=-\frac{\partial\log p_{t,\theta}(x_{t})}{\partial x_{t}}^{T}\frac{\partial v_{\theta}(x_{t},t)}{\partial x_{t}}-\frac{\partial}{\partial x_{t}}Tr\left(\frac{\partial v_{\theta}(x_{t},t)}{\partial x_{t}}\right)$ (19)
@@
$Tr\left(\frac{\partial v_{\theta}(x_{t},t)}{\partial x_{t}}\right)\,$ (21)
@@ 4 Experiments
Motivated by the improvements on the toy model, we now turn to more challenging problems, namely the experiments done in Klein et al. (2023b) and Klein & Noé (2024). To this end, we use the same architecture, a CNF with an EGNN (E(n) Equivariant Graph Neural Network) (Satorras et al., 2021; Garcia Satorras et al., 2021) for the vector field. The architecture in Klein & Noé (2024) is mostly the same as in Klein et al. (2023b), with the main difference of the encoding of the atoms. While the model in Klein et al. (2023b) treats the same atom type as indistinguishable, the model in Klein & Noé (2024) encodes nearly all atoms differently. Only hydrogens bound to the same atom are treated as indistinguishable. For more details on the model architecture see Section˜A.2. While we maintain the architecture and the initial training process with Flow Matching, we change the later training procedure to include Path Gradients. As in Klein et al. (2023b) we investigate optimal transport Flow Matching (OT FM), – and for the second set of experiments on LJ13 – equivariant optimal transport Flow Matching (EQ OT FM), and naive Flow Matching (standard FM).
We first evaluate the models on two many particle system with pair-wise Lennard-Jones interactions with 13 and 55 particles. In addition we investigate Alanine dipeptide (AD2) at $T=300K$ both with a classical and a semi-empirical force field (XTB). In contrast to (Klein et al., 2023b; Klein & Noé, 2024), we do not bias the training data towards the positive $\varphi$ state. The bias was originally introduced to make learning problem simpler, as the unlikely state is more prominent (Klein et al., 2023b). This, however, skews the evaluation, since the metrics, namely ESS and Negative log likelihood (NLL), are based on the target distribution, not the biased one. Moreover, this assumes system-specific knowledge of slow-varying degrees of freedom, information that is typically unavailable for unknown systems. Finally, we also investigate the 2AA dataset consisting of dipeptides simulated at $T=310K$ with a classical force field (Klein et al., 2023a), to evaluate the influence of PG in a transferability. For more details on the datasets see Section˜A.3. We published code for replicating our experiments 222github.com/lenz3000/path-grads-after-fm. In total, we investigate three scenarios:
@@
  • Second, we apply Path Gradients on the FM-trained models using unlimited resources to maximize performance after training with Standard FM, OT FM and EQ OT FM as done in Klein et al. (2023b), see Section˜4.2. In the second case, we additionally examine the ODE integration length to understand how PG fine-tuning influences the model.
@@ 4.1 Fine-tuning using the same limited resources
We observe that fine-tuning with Path Gradients improves the ESS metrics, i.e. ESSp and ESSq, across most datasets and models, despite their high variance. The hybrid approach reliably improves the NLL for all models and datasets, but one. The only exception is the standard Flow (Klein et al., 2023b) on AD2 with XTB, where fine-tuning with FM still performs better on average. Examining the ESS, we find that both approaches fail to adequately fit the target density. This supports the hypothesis that Path Gradients are only effective for fine-tuning, when the flow already provides a reasonably good approximation of the target. Notably, on LJ55, our hybrid approach nearly triples the ESS, and on AD2 with XTB, it doubles it – in the same training time.
@@ 4.3 Transferability on dipeptides
Finally, we also investigate whether fine-tuning with Path Gradients improves performance in the transferable setting. To this end we fine-tune Transferable Boltzmann Generators (TBG) (Klein & Noé, 2024), which were trained on 200 different dipeptides and evaluate on 16 unseen ones, like in Klein & Noé (2024). The experiments show that fine-tuning with path gradients improves the NLL and energies for all evaluated unseen test dipeptides. Further, on average the ESSq is improved by around 23% to an efficiency of $9.79\%$. For more details see Section˜A.10.
@@ A.8 Comparison with maximum likelihood training
We here present comparisons with maximum likelihood training as a baseline. To this end we compare to Garcia Satorras et al. (2021), which is nearly the same architecture as in Klein et al. (2023b), but the model is trained via maximum likelihood training instead of flow matching. The results are shown in Table˜4.
@@ A.11.2 Variance of the estimators
In the case of perfect approximation, i.e. $q_{\theta}(x)=p(x)\forall x\in\mathcal{X}$, the following statements about the ML estimator and the Path Gradient estimators are known. The gradients for path gradient estimators are deterministically 0, i.e. $\mathbb{E}[\mathcal{G}_{PG}]=0,\mathrm{Var}[\mathcal{G}_{PG}]=0$, while for ML the variance is generically nonzero $\mathbb{E}[\mathcal{G}_{ML}]=0,\mathrm{Var}[\mathcal{G}_{ML}]=\frac{1}{N}\mathcal{I}_{\theta}$. Where $\mathcal{I}_{\theta}$ is the Fisher Information Matrix of $p_{0,\theta}$ [2].

### v3

@@ 1 Introduction
  • We investigate Path Gradients’ impact on properties trained during Flow Matching, namely flow trajectory length and the MSE Flow Matching loss, and show that these are mostly unaffected by the fine-tuning.
@@ 2.1 Boltzmann Generators
Boltzmann Generators (Noé et al., 2019) aim to asymptotically generate unbiased samples from the equilibrium Boltzmann distribution
@@ 2.4 Path Gradients
Roeder et al. (2017) observed, that the martingale score term vanishes in expectation, but has a non-zero variance. Vaitl et al. (2022a) showed that its variance is $\frac{1}{N}I_{\theta}$ for a batch of size $N$ and the Fisher Information $I_{\theta}$ of the variational distribution $q_{\theta}$. The low variance close to the optimum allows Path Gradient estimators to "stick-the-landing" (Roeder et al., 2017), i.e. having zero variance at the optimum, making it an ideal tool for fine-tuning. Vaitl et al. (2024) showed that Path Gradient estimators incorporate additional information about the derivative of both target and variational distribution, opposed to the reparametrized gradient gradient estimator (Kingma & Welling, 2014; Mohamed et al., 2020) (see Section˜A.11.1 for a detailed explanation). They hypothesized that this leads to less overfitting and a generally better fit to the target density (see Vaitl (2024)). These theoretical aspects make Path Gradients a promising suitor for fine-tuning Boltzmann Generators.
@@
While the naive calculation of Path Gradients requires significantly more compute and memory than Flow Matching, we take several steps to obtain constant memory and speed ups in the next section. As we show experimentally, fine-tuning with Path Gradient for a few epochs significantly improves the performance for Continuous Normalizing Flows compared to only using Flow Matching.
@@ 3 Path Gradients and Flow Matching
$\frac{d}{dt}\frac{\partial\log p_{t,\theta}(x_{t})}{\partial x_{t}}=-\frac{\partial\log p_{t,\theta}(x_{t})}{\partial x_{t}}^{T}\frac{\partial v_{\theta}(x_{t},t)}{\partial x_{t}}-\frac{\partial}{\partial x_{t}}\text{Tr}\left(\frac{\partial v_{\theta}(x_{t},t)}{\partial x_{t}}\right)$ (19)
@@
$\text{Tr}\left(\frac{\partial v_{\theta}(x_{t},t)}{\partial x_{t}}\right)\,$ (21)
@@ 4 Experiments
Motivated by the improvements on the toy model, we now turn to more challenging problems, namely the experiments done in Klein et al. (2023b) and Klein & Noé (2024). To this end, we use the same architecture, a CNF with an EGNN (E(n) Equivariant Graph Neural Network) (Satorras et al., 2021; Garcia Satorras et al., 2021) for the vector field. The architecture in Klein & Noé (2024) is mostly the same as in Klein et al. (2023b), with the main difference of the encoding of the atoms. While the model in Klein et al. (2023b) treats the same atom type as indistinguishable, the model in Klein & Noé (2024) encodes nearly all atoms differently. Only hydrogens bound to the same atom are treated as indistinguishable. For more details on the model architecture see Section˜A.2. While we maintain the architecture and the initial training process with Flow Matching, we change the later training procedure to include Path Gradients. As in Klein et al. (2023b), we investigate optimal transport Flow Matching (OT FM), – and for the second set of experiments on LJ13 – equivariant optimal transport Flow Matching (EQ OT FM), and naive Flow Matching (standard FM).
We first evaluate the models on two many particle system with pair-wise Lennard-Jones interactions with 13 and 55 particles. In addition we investigate Alanine dipeptide (AD2) at $T=300K$ both with a classical and a semi-empirical force field (XTB). In contrast to (Klein et al., 2023b; Klein & Noé, 2024), we do not bias the training data towards the positive $\varphi$ state. The bias was originally introduced to make learning problem simpler, as the unlikely state is more prominent (Klein et al., 2023b). This, however, skews the evaluation, since the metrics, namely ESS and Negative log likelihood (NLL), are based on the target distribution, not the biased one. Moreover, it assumes system-specific knowledge of slow-varying degrees of freedom, information that is typically unavailable for unknown systems. Finally, we also investigate the 2AA dataset consisting of dipeptides simulated at $T=310K$ with a classical force field (Klein et al., 2023a), to evaluate the influence of PG in a transferability. For more details on the datasets see Section˜A.3. We published code for replicating our experiments 222github.com/lenz3000/path-grads-after-fm. In total, we investigate three scenarios:
@@
  • Second, we apply Path Gradients on the FM-trained models using unlimited resources to maximize performance after training with Standard FM, OT FM and EQ OT FM as done in Klein et al. (2023b), see Section˜4.2. In this second case, we additionally examine the ODE integration length to understand how PG fine-tuning influences the model.
@@ 4.1 Fine-tuning using the same limited resources
We observe that fine-tuning with Path Gradients improves the ESS metrics, i.e. ESSp and ESSq, across most datasets and models, despite their high variance. The hybrid approach reliably improves the NLL for all models and datasets, but one. The only exception is the standard Flow (Klein et al., 2023b) on AD2 with XTB, where fine-tuning with FM still performs better on average. Examining the ESS, we find that both approaches fail to adequately fit the target density. This supports the hypothesis that Path Gradients are only effective for fine-tuning when the flow already provides a reasonably good approximation of the target. Notably, on LJ55, our hybrid approach nearly triples the ESS, and on AD2 with XTB, it doubles it – in the same training time.
@@ 4.3 Transferability on dipeptides
Finally, we also investigate whether fine-tuning with Path Gradients improves performance in the transferable setting. To this end we fine-tune Transferable Boltzmann Generators (TBG) (Klein & Noé, 2024), which were trained on 200 different dipeptides and evaluate on 16 unseen ones, like in Klein & Noé (2024). The experiments show that fine-tuning with path gradients improves the NLL and energies for all evaluated unseen test dipeptides. On average, we observe a relative improvement of around 23% in the ESSq, reaching an efficiency of $9.79\%$. For more details see Section˜A.10.
@@ A.8 Comparison with maximum likelihood training
We here present comparisons with maximum likelihood training as a baseline on LJ13. To this end we compare to Garcia Satorras et al. (2021), which is nearly the same architecture as in Klein et al. (2023b), but the model is trained via maximum likelihood training instead of flow matching. The results are shown in Table˜4.
@@ A.11.2 Variance of the estimators
In the case of perfect approximation, i.e. $q_{\theta}(x)=p(x)\forall x\in\mathcal{X}$, the following statements about the ML estimator and the Path Gradient estimators are known. The gradients for path gradient estimators are deterministically 0, i.e. $\mathbb{E}[\mathcal{G}_{PG}]=0,\mathrm{Var}[\mathcal{G}_{PG}]=0$, while for ML the variance is generically nonzero $\mathbb{E}[\mathcal{G}_{ML}]=0,\mathrm{Var}[\mathcal{G}_{ML}]=\frac{1}{N}\mathcal{I}_{\theta}$. Where $\mathcal{I}_{\theta}$ is the Fisher Information Matrix of $p_{0,\theta}$ (Vaitl et al., 2022a).